# Video-based Automatic Lameness Detection of Dairy Cows using Pose Estimation and Multiple Locomotion Traits

### v1

@@ 5.4 Feature importance
Multiple studies investigated the relationship between individual locomotion traits and locomotion scores (flower_effect_2006, ; borderas2008effect, ; chapinal2009using, ; schlageter-tello_relation_2015, ). They found that, when scored individually, the traits arched back, asymmetric gait, head bobbing, reluctance to bear weight and tracking-up were highly correlated with the locomotion score. The features selected in this study were designed to measure the same traits. The arched back was measured by the Back Curvature Measurement (HBA), the asymmetric gait by the Stride Length (STL) difference between left and right limbs, the head bobbing by the Head Bobbing Amplitude (HBA), the reluctance to bear weight by the Stance Duration (STD) and Swing Duration (SWD), and the tracking up was measured by the Tracking distance (TRK).
@@
The BPM, HBA, and TRK features returned the highest scores in the permutation importance test. BPM and HBA displayed a clear demarcation between the normal and lame classes in Figure 6. As reported by flower_effect_2006 borderas2008effect and chapinal2009using , it suggests that the back posture, head bobbing, and tracking-up are, for human observers, easier to recognize than an asymmetric gait (e.g. stride length). The tracking distance on the left side (TRK-L) had a higher importance than the one on the right side (TRK-R). This could indicate that, in our dataset, there were more cows tracking-up on the left than on the right side.
@@
Both for the Stance Duration (STD) and the Swing Duration (SWD) on the hind legs (Fig. 6), one can see a clear difference in the duration of the stance/swing phases between the classes, whereas classes differences are less obvious on the front legs. This could be explained by the fact that lameness happens more often on the hind legs (poursaberi_real-time_2010, ; flower_effect_2006, ). Including SWD as a feature increased the classification performance, even though SWD had the lowest importance score. In contrast, STD had a larger importance score than SWD, but adding the STD feature to the input of the classifier led to a small decrease in accuracy and F1-score. This could indicate multi-collinearity with other features.
@@
The STL features had the second lowest importance score and the class separation was harder to distinguish in Figure 6. Interestingly, the F1-score, sensitivity, and specificity were higher when the STL features were included. This suggests that the stride length can be informative when used in combination with other features. It is worth noting that if the cows have bilateral lameness, i.e., are lame on left and right limbs, then the stride length would show little to no difference (blackie_associations_2013, ).
@@
Overall, combining multiple locomotion traits led to a better classification performance than using a single trait. Using a combination of 3 and 6 traits led to the best accuracy and F1-scores on the SVM classifier with a radial kernel. Even though additional traits could be extracted from the keypoint trajectories, it is unknown whether they would lead to significant improvements in the gait classification. Our recommendation would be to include at least the following locomotion traits in an automatic lameness detection system: back posture, head bobbing, and tracking distance, as they demonstrated good overall classification metrics, and these features have been shown to be highly correlated with the locomotion scores (flower_effect_2006, ; borderas2008effect, ; chapinal2009using, ).
@@ 5.5 Comparison with related work
Directly comparing the performance of our lameness classifiers against related work is not straightforward, because even though the task at hand (i.e., detecting lameness from videos) is the same, there is a large variation in the material, methods, and evaluations used in papers that address it. Furthermore, a comprehensive literature review is out of the scope of this paper, and we refer the reader to nejati2023technology for an overview of past and current advances in bovine gait analysis. We will here compare our results and contrast our findings with previous work that we deem directly related to ours.
@@
The Back Posture Measurement (BPM) was first introduced by poursaberi_real-time_2010 and curvature of the back has since then been used in numerous studies poursaberi_real-time_2010 ; viazzi_analysis_2012 ; van_hertem_automatic_2014 ; viazzi_comparison_2014 ; van_hertem_implementation_2018 ; jiang_dairy_2022 ; zheng_cows_2023 ; zhao_automatic_2023 ; barney2023deep . The BPM is commonly measured during the supporting phase of the hind hooves, and not during the supporting phase of the front hooves because lameness is more common on the hind hooves than on the front ones. However, this practice could lead to front lameness cases being systematically missed by the algorithm. To prevent this, we computed the BPM based on the supporting phase of the four legs. When using BPM as a single locomotion trait, the accuracy of lameness classification ranged from 76% (viazzi_analysis_2012, ) to 96% (jiang_dairy_2022, ). When only including the BPM trait in our SVM-R classifier, we reached an accuracy of 76.6%, which is in line with the literature.
@@
The work presented in zhao_automatic_2023 is perhaps the most closely related to this study. In zhao_automatic_2023 , the authors used a combination of traditional and deep-learning-based computer vision to develop a lameness detection system. They used DeepLabCut mathis_deeplabcut_2018 , a deep-learning model that was trained to track the location of the hoofs and the head in videos of walking cows without physical markers. However, a pixel-level background subtraction method was used for extracting the outline of the spine, which might not be robust to varying backgrounds and light. The videos where the keypoint predictions were too erroneous were manually discarded. In total, they used 212 videos of walking cows, where cows that were given a score of 1 or 2 were classified as normal, and a score of 3 or 4 as lame. The back curvature was computed from the outline of the spine, and the keypoints on the hooves and on the neck were used to extract the following features: head bobbing, stride length asymmetry, tracking up, landing speed, supporting phase asymmetry, and moving speed. The feature selection was performed as follows: a Chi-square test was run on the whole dataset. The test revealed that back posture measurement and head bobbing were the most important features. Several classifiers were trained with the back curvature and head bobbing, and the logistic regression classifier returned the best results, with a classification accuracy of 87.3%. They reported that no other combination of features performed better than back curvature and head bobbing. In contrast, we found that adding tracking-up to the other two features led to better results on our dataset. This could mean that, in their dataset, lame subjects were not tracking up. Another explanation could be that with increasing the number of traits, the complexity of the data increases, and a non-linear classifier, such as SVM-R, would be needed.
@@
In barney2023deep , a fully automated multi-cow lameness detection system was developed. They used a Mask-R-CNN, a deep-learning model, to simultaneously perform object-detection of the cows, and pose estimation of 7 keypoints located on the back neck and head. In total, they used 250 videos of 10 different cows. The keypoints were used to extract the back curvature and head position locomotion traits. Each locomotion trait was extracted per video frame and aggregated per video into statistical features such as the mean, median, standard deviation, min, and max values. They trained the CatBoost gradient boosting classifier and achieved a 98% accuracy on binary lameness detection, and 94% accuracy on a 4-point scale lameness scoring. In our work, although we included four more locomotion traits, we only aggregated the values into the median value of the video. In light of the excellent performance of their classifiers, a promising direction for extending our work would then be to extract more statistical features from the locomotion traits, such as mean, standard deviation, and min and max values, to further improve our classification performance.

### v2

@@ 5.4 Feature importance
Multiple studies investigated the relationship between individual locomotion traits and locomotion scores [28, 42, 43, 25]. They found that, when scored individually, the traits arched back, asymmetric gait, head bobbing, reluctance to bear weight and tracking-up were highly correlated with the locomotion score. The features selected in this study were designed to measure the same traits. The arched back was measured by the Back Curvature Measurement (HBA), the asymmetric gait by the Stride Length (STL) difference between left and right limbs, the head bobbing by the Head Bobbing Amplitude (HBA), the reluctance to bear weight by the Stance Duration (STD) and Swing Duration (SWD), and the tracking up was measured by the Tracking distance (TRK).
@@
The BPM, HBA, and TRK features returned the highest scores in the permutation importance test. BPM and HBA displayed a clear demarcation between the normal and lame classes in Figure 6. As reported by [28][42] and [43], it suggests that the back posture, head bobbing, and tracking-up are, for human observers, easier to recognize than an asymmetric gait (e.g. stride length). The tracking distance on the left side (TRK-L) had a higher importance than the one on the right side (TRK-R). This could indicate that, in our dataset, there were more cows tracking-up on the left than on the right side.
@@
Both for the Stance Duration (STD) and the Swing Duration (SWD) on the hind legs (Fig. 6), one can see a clear difference in the duration of the stance/swing phases between the classes, whereas classes differences are less obvious on the front legs. This could be explained by the fact that lameness happens more often on the hind legs [6, 28]. Including SWD as a feature increased the classification performance, even though SWD had the lowest importance score. In contrast, STD had a larger importance score than SWD, but adding the STD feature to the input of the classifier led to a small decrease in accuracy and F1-score. This could indicate multi-collinearity with other features.
@@
The STL features had the second lowest importance score and the class separation was harder to distinguish in Figure 6. Interestingly, the F1-score, sensitivity, and specificity were higher when the STL features were included. This suggests that the stride length can be informative when used in combination with other features. It is worth noting that if the cows have bilateral lameness, i.e., are lame on left and right limbs, then the stride length would show little to no difference [9].
@@
Overall, combining multiple locomotion traits led to a better classification performance than using a single trait. Using a combination of 3 and 6 traits led to the best accuracy and F1-scores on the SVM classifier with a radial kernel. Even though additional traits could be extracted from the keypoint trajectories, it is unknown whether they would lead to significant improvements in the gait classification. Our recommendation would be to include at least the following locomotion traits in an automatic lameness detection system: back posture, head bobbing, and tracking distance, as they demonstrated good overall classification metrics, and these features have been shown to be highly correlated with the locomotion scores [28, 42, 43].
@@ 5.5 Comparison with related work
Directly comparing the performance of our lameness classifiers against related work is not straightforward, because even though the task at hand (i.e., detecting lameness from videos) is the same, there is a large variation in the material, methods, and evaluations used in papers that address it. Furthermore, a comprehensive literature review is out of the scope of this paper, and we refer the reader to [44] for an overview of past and current advances in bovine gait analysis. We will here compare our results and contrast our findings with previous work that we deem directly related to ours.
@@
The Back Posture Measurement (BPM) was first introduced by [6] and curvature of the back has since then been used in numerous studies [6, 19, 20, 21, 22, 13, 7, 8, 17]. The BPM is commonly measured during the supporting phase of the hind hooves, and not during the supporting phase of the front hooves because lameness is more common on the hind hooves than on the front ones. However, this practice could lead to the algorithm systematically missing front lameness cases. To prevent this, we computed the BPM based on the supporting phase of the four legs. When using BPM as a single locomotion trait, the accuracy of lameness classification ranged from 76% [19] to 96% [13]. When only including the BPM feature in our SVM-R classifier, we reached an accuracy of 76.6%, which is in line with the literature.
The work presented in [8] is perhaps the most closely related to this study. In [8], the authors used a combination of traditional and deep-learning-based computer vision to develop a lameness detection system. They used DeepLabCut [15], a deep-learning model trained to track the location of the hoofs and the head in videos of walking cows without physical markers. The pose estimation model achieved a percentage of correct keypoints (PCK) of 92% and the videos where the keypoint predictions were too erroneous were manually discarded. The outline of the cow’s spine was obtained using a background subtraction method at the pixel level. However, this method may not be effective in handling varying backgrounds and lighting conditions. On the other hand, we used pose estimation to track all the keypoints on the cow’s body, including those on the spine. As a result, our proposed method is more robust in dealing with such changes. Additionally, we automatically corrected erroneous keypoint trajectories with our keypoint-correction and step-correction approaches, which removed the need for manual validation.
@@
In total, they used 212 videos of walking cows, where cows with a score of 1 or 2 were classified as normal, and a score of 3 or 4 as lame. The back curvature was computed from the outline of the spine, and the keypoints on the hooves and on the neck were used to extract the following features: head bobbing, stride length asymmetry, tracking up, landing speed, supporting phase asymmetry, and moving speed. The back curvature and head bobbing were computed when the hind hooves were in contact with the floor and, therefore, didn’t account for lameness on the front limbs. In contrast, we computed the back curvature over all four legs, and the head bobbing amplitude was extracted from the entire trajectory. The feature selection was performed as follows: a Chi-square test was run on the whole dataset. The test revealed that back posture measurement and head bobbing were the most important features. In contrast, we found that adding tracking-up to the other two features led to better results on our dataset. This could mean that, in their dataset, lame subjects were not tracking up. Another explanation could be that as the number of traits increases, the complexity of the data increases, and a non-linear classifier, such as SVM-R, would be needed. Several classifiers were trained with the back curvature and head bobbing, and the logistic regression classifier returned the best results, with a classification accuracy of 87.3%. In comparison, our accuracy of 80.07% may seem modest. However, it’s important to note that there are several differences in the data and experimental setup that could have impacted the results. The datasets and methods used in the studies differed, making a direct comparison challenging. In our study, most of the locomotion scores were on the lower end of the scale, and the small number of severely lame cows may have made it harder for the classifier to distinguish between lame and non-lame cows. Additionally, the quality of the ground truth data might have affected the performance of our classifiers. We also followed best machine-learning practices by conducting feature selection only on the training sets and by separating individual cows in the training and validation sets to prevent data leakage. Failing to do so can artificially inflate performance results and lead to overly optimistic conclusions [45]. Overall, we would expect our method to yield similar results if it were run on their dataset, with the advantage of a fully automatic pipeline that doesn’t require manual validation of the keypoint trajectories, and a pose estimation method robust to light conditions and occlusions.
@@
In [17], a fully automated multi-cow lameness detection system was developed. They used a Mask-R-CNN, a deep-learning model, to simultaneously perform object-detection of the cows, and pose estimation of 7 keypoints on the back neck and head. In total, they used 250 videos of 10 different cows. The keypoints were used to extract the back curvature and head position locomotion traits. Each locomotion trait was extracted per video frame and aggregated per video into statistical features such as the mean, median, standard deviation, min, and max values. They trained the CatBoost gradient boosting classifier and achieved a 98% accuracy on binary lameness detection and 94% accuracy on a 4-point scale lameness scoring. In our work, although we included four more locomotion traits, we only aggregated the values into the median value of the video. In light of their classifiers’ excellent performance, a promising direction for extending our work would be to extract more statistical features from the locomotion traits, such as mean, standard deviation, and min and max values, to improve our classification performance further.